OpenStudy (anonymous):

Katalene wants to make a quilt in the shape of a golden rectangle. She only has enough binding material for a quilt with a perimeter of 29 feet. What should the dimension og the quilt be? ... A. 5.44 ft by 9.06 ft ... B) 5.54 ft by 8.96 ft ... C) 5.74 ft by 8.76 ft ... D) 5.64 ft by 8.86 ft

OpenStudy (anonymous):

Hint: the sides of a golden rectangle are in the ratio \(1:\dfrac{1+\sqrt{5}}{2}\).

Directrix (directrix):

The Golden Ratio is a "rich" topic in mathematics. After you complete this problem, you can read about it at this link: http://www.mathsisfun.com/numbers/golden-ratio.html

Directrix (directrix):

Take each of your options, place the longer dimension value over the smaller, divide, and determine which option comes closest to 1.618 which is an approximated Golden Ratio. For example, consider option D. 8.86/5.64 = 1.57092198582. That's off so try another option. Note: I entered 8.86/5.64= in the Google search box and the 1.57092198582 popped up in the Google calculator. @xayemeehx